OpenStudy (anonymous):

write the equation of the line that is parallel to the line 4x - 3 = -12 and passes through the point (-3,4)

OpenStudy (anonymous):

please help

rishavraj (rishavraj):

if two lines are parallel then their slopes are equal...

OpenStudy (anonymous):

then?

rishavraj (rishavraj):

once u get the slope (m) \[\frac{ y_2 - y_1 }{ x_2 - x_1 } = m\]

OpenStudy (anonymous):

i forgot the answers *slaps myself*

rishavraj (rishavraj):

forgot the answers????wht do u mean???? huh just plug hmmmm y_1 = 4 aand x_1 = -3 and m

OpenStudy (anonymous):

Here are the answers A. y = 4\3x + 8 B. y = 4\3x + 3 C. y = -3\4x + 8 D. y = -3\4x + 3

rishavraj (rishavraj):

4x - 3 = -12 ?????o_O its 4x - 3y = -12 -_- did u get the slope of required line???

rishavraj (rishavraj):

first of all write it in the format of y = mx + c so 4x - 3y = -12 y = (4/3)x + 4 so slope i.e m = 4/3 hence required line is parallel so slopes are equal i.e 4/3

rishavraj (rishavraj):

@rikkibracamonte1999

OpenStudy (anonymous):

i am going with option C

rishavraj (rishavraj):

cool as u wish ....bt thts wrong.....

OpenStudy (anonymous):

really? :(

OpenStudy (anonymous):

i suck at math so i come here for help

rishavraj (rishavraj):

yup ..... u got the slope i.i 4/3 now just solve \[\frac{ y - 4 }{ x + 3 } = \frac{ 4 }{ 3 } \]

OpenStudy (anonymous):

so its either a or b?

rishavraj (rishavraj):

yup it is one of them.....:)

OpenStudy (anonymous):

my gut is telling me its a

rishavraj (rishavraj):

yup it is........

OpenStudy (anonymous):

 sorry this isn't my best drawing

OpenStudy (anonymous):

i tried :/

OpenStudy (anonymous):

A. y = 0 B. y = -2 C. y = 3 D. y = 3x ^ for the drawing + question

rishavraj (rishavraj):

@rikkibracamonte1999 upload a screenshot of the figure,,,,,